# Underwater Image Enhancement by Diffusion Model with Customized CLIP-Classifier

### v1

@@ IV-A Benchmarks and Evaluation Metrics
Prior Knowledge. It is well known that real reference images are difficult to obtain for underwater image enhancement tasks. We adopt the synthetic dataset strategy, using color transfer to degrade the in-air natural images of INaturalist 2021 (mini) [34] to the underwater domain to yield paired datasets (see in Fig. 2 (a)). INaturalist 2021 (mini) has 10,000 categories with 50 images in each category, totalling 500k images. The image-to-image diffusion model is trained from scratch on this dataset to learn prior knowledge from the real underwater degradation domain to the in-air natural domain.
Dataset. The direct finetuning or continued training of a large pretrained image-to-image diffusion model with limited data may cause overfitting and catastrophic forgetting [21]. The SUIM-E [11] has only 1635 images, while UIEB [17] has only 890 images, which is significantly smaller than the INaturalist dataset used to pre-train the diffusion model. Therefore, we utilize the pre-divided SUIM-E-Train set, which consists of 1525 paired underwater images, and 800 paired images randomly selected from UIEB as the training set, for fine-tuning our model. The test set from SUIM-E and the remaining 90 images from UIEB are combined to form a test set of 200 images, called T200.
We also conduce tests on the Challenging set of UIEB, for which contains 60 tough cases for enhancement without references, called C60. In addition, we further verify the generalisation and robustness of our model on the following commonly used underwater image datasets, which do not have reference images. SQUID [51] contains 57 images taken in different locations, enabling a rigorous evaluation of restoration algorithms on in-air natural images. Color-Checker7 [52] covers 7 underwater images taken with the camera from different angles in a shallow swimming pool, which is also used to evaluate the accuracy of the proposed CLIP-UIE for color correction.
Evaluation Metrics. Since not all datasets have reference images, we divide five commonly used evaluation metrics into two groups for comprehensive quantitative evaluation. For full-reference evaluation metrics, we use two metrics: Peak Signal to Noise Ratio (PSNR) which measures the ratio of the peak signal to the average energy, and Structural Similarity (SSIM[53] which compares the target and ground-truth images in terms of brightness, contrast, and structure. For non-reference evaluation metrics, we utilize three metrics: Underwater Image Quality Measure (UIQM[54] which considers the underwater contrast, colorfulness, and sharpness on HSV space, Underwater Color Image Quality Evaluation (UCIQE[55] which is based on chroma, contrast, and saturation to quantify the nonuniform color cast, blurring, and low-contrast for the degraded underwater images, and Cumulative Probability of Blur Detection (CPBD[56] which utilizes a probabilistic model to estimate the probability of detecting blur at each edge in the image.
@@ IV-B Implementation Details
We implement our method with the PyTorch platform. For training, we initialize all the weights of each layer with Xavier initialization [57] and employ the Adam optimizer with a learning rate of 3e-6 with the scheduler of linear decay to zero. Moreover, batch size is set to 8, and the time step of the diffusion model is set to 2000 (namely T=2000). To ensure a fair comparison, images for all compared methods were resized to 256×256 during both training and testing. Additionally, our data augmentation techniques include random rotation and random horizontal flips. The experiments are conducted on a consistent hardware setup comprising a single NVIDIA RTX 3090 GPU, a 2.3GHz Intel Xeon processor, 128GB RAM, and the Ubuntu 18.04 operating system platform. The pre-trained model we used has been trained from scratch for 300 hours on INaturalist 2021 (mini).
@@ IV-C Compared Methods
We compare our method CLIP-UIE with seven state-of-the-art UIE approaches, including two typical traditional methods (UDCP [8] and ULAP [10]), and five deep learning-based methods (Ucolor [1], TCTL-Net [30], UIEC^2-Net [12], UDAformer [15], DM_underwater [41]). The aforementioned two UIE methods are the most representative physical model-free methods that have been widely adopted in recent years. The deep learning-based approaches compared here have achieved state-of-the-art (SOTA) performance over the past years. Among them, the Ucolor [1] is based on medium transmission-guided multi-color space embedding, the TCTL-Net [30] is based on template-free color transfer, UIEC^2-Net [12] is based on RGB and HSV Color Space, UDAformer [15] employs dual attention transformer, and DM_underwater [41] is the first UIE method based on diffusion model. To ensure a fair and rigorous comparison, we utilized the same training set from the SUIM-E and UIEB dataset to fine-tune or continued train all the deep learning-based methods. Since the Ucolor [1] training phase requires special image sizes, we only use 256x256 image sizes for the test phase. Other comparison methods use 256x256 image size in both training and testing phases. In addition, all the comparison methods utilize the source codes released by the respective authors to produce their results and strictly adhere to the identical experimental settings across all evaluations.
@@
IV-D Comparison on Visual Quality of Enhancement
The quantitative comparisons on the T200, C60, SQUID and Color-Check7 datasets are reported in Table I. We mainly use PSNR and SSIM as our quantitative indices for full-reference T200 datasets, and UIQM, UCIQE and CPBD for non-reference C60, SQUID and Color-Checker7 datasets. The results in Tab. I show that our CLIP-UIE outperforms the other compared methods, and achieve state-of-the-art performance in terms of image quality evaluation metrics on UIE task, particularly in UIQM and UCIQE metrics, with an improvement of 0.1 and 0.01, respectively. This verifies that our method can produce more natural enhanced images, which better corrects degraded colors and maps from underwater degraded domains to in-air natural domains. To better validate the superiority of our methods, Figure 6 shows the visual results of different comparison methods on T200 dataset. The traditional methods like UDCP and ULAP, usually cause color distortions and damage image detail (see the examples in Fig. 6 (b, c)), albeit with high non-reference metrics, which also occurs in other datasets. The deep learning-based methods can produce pleasing visual enhancement due to the direct use of reference images as constraints during the training phrase. However, this limits the enhanced image to a human-selected reference domain, and some methods have very similar enhancement results and do not outperform the reference (e.g. (d - h) in the second and fourth rows of Fig. 6). On the contrary, our CLIP-UIE introduces the prior knowledge of in-air natural domain through CLIP-Classifier, which can produce more natural images, breaking through the limitation imposed by reference domain to a certain extent, as shown in Fig. 6 (i).
To further verify the effectiveness and robustness of our method, we conducted experiments on C60, SQUID and Color-Checker7 datasets. (1) The visual comparisons with other methods on the C60 dataset are shown in the Fig. 7. The underwater images suffer from lower contrast and more severe color casts (see the examples in the first row of Fig. 7), which raise great challenges for the enhancement methods. By observation, the deep learning-based methods show better robustness and generalisation than the traditional methods. As a comparison, our method has better performance in color correction and detail restoration. Specially, our method improves 11% in UIQM metrics. (2) On the SQUID dataset, our CLIP-UIE outperforms all other deep learning-based methods and achieves optimal scores in terms of underwater image quality non-reference evaluation metrics including UIQM and UCIQE, with 0.424 and 0.575, respectively. (see in Table I). As shown in Figure 8, our method performs more natural visual enhancement for underwater images taken at different depths, which verifies the robustness of the proposed CLIP-UIE. This phenomenon can be attributed to the conditional generation is jointly controlled by the source image $y_{1}$ and CLIP-Classifier $y_{2}$ (the process is shown in Fig. 3), thereby the multi-conditions guidance enhances generation quality. (3) We further conduce comparative tests on the Color-Checker7 dataset to evaluate the generalization and robustness of our model. Figure 9 visually illustrates the difference between output images of CLIP-UIE and other methods. In the first row of Fig. 9, the difference lies in the correction of skin color, where the image generated by CLIP-UIE is closer to the real colors. In the second row of Fig. 9, the ULAP [10] produce the enhanced image with good background lighting, but there is still a color deviation due to the introduction of excessive red components. Comparing to deep learning-based methods, the results produced by our method are brighter and more visually appealing.
@@ IV-D Comparison on Visual Quality of Enhancement
IV-E Comparison on Color Correction Performance
In order to further illustrate the performance of the proposed CLIP-UIE for color restoration of underwater images, we conduced detailed color dissimilarity comparisons on the Color-Checker7 [52] dataset, which is photographed in a swimming pool with divers holding standard color cards. Following Ancuti et al. [10], we use CIEDE2000 [52] to measure the relative differences between the corresponding color patches of the Macbeth color checker [58] and the enhancement results of these comparison state-of-the-art methods, making it easy to understand that a standard color card is appended to the upper left corner of the first column of Fig. 9. The results, as presented in Tab. II, indicates that improve the performance of UIE methods in recovering ground-truth colors, but not very noticeably, and the enhancement result of the image T8000 is even worse than the input. We believe that the image encoder $\Phi(\cdot)$ compresses images into high-dimensional features, and CLIP-Classifier pays more attention to global features during generation process. As shown in the first row of Fig. 9, the result of our method presents better brightness and visual perception, but does not achieve a significant improvement in color correction.
@@ IV-F Ablation Study
Effectiveness of CLIP-Classifier. To Investigate the impact of the CLIP-Classifier on our method, we fine-tune our conditional diffusion model on the same training data without CLIP-Classifier (w/o-CLIP). The scores of the main evaluation metrics on T200 datasets are given in Tab. III. The results indicate that CLIP-Classifier guides the diffusion model toward the in-air natural domain, leading to a decrease in the full-reference metrics and an increase in the non-reference metrics. Besides, we use CLIP-Classifier to evaluate the similarity (CLIP score) between the in-air natural domain and enhanced results, and the corresponding point-box plot and kernel density estimation (KDE) curve are shown in Fig. 10. The distributions of the w/o-CLIP and the Reference are very similar (see KDE in Fig. 10), which indicates that its enhancement performance is constrained by the reference domain. On the contrary, due to the introduction of the prior knowledge of in-air natural domain through CLIP-Classifier, our CLIP-UIE breaks through this limitation imposed by reference domain to a certain extent, keeping the distribution of the CLIP-UIE far away from the reference domain and close to the in-air natural domain and reducing subjective preferences.
Visual comparison illustrated in Figure 11 can also significantly support the above points. Compared with its ablated versions, the proposed CLIP-UIE achieves more natural enhancement results, with the guidance of CLIP-Classifier that inherits prior knowledge of the in-air natural domain. And without real-world underwater reference images, the combination of the reference and in-air natural domains drives the enhancement results closer to real-world natural images (see Fig. 3 and Fig. 11 (c, d)).
@@
Fine-Tuning Strategy. In addition to the observation provided in Section III-C, to further validate the new fine-tuning strategy, we provide the quantitative comparison in Tab. IV. The model $t_{m}$ represents the fact that we only fine-tune the model pretrained on INaturalist [34] in the range $t\in(0,t_{m})$ (the training time step $n\in(0,2000\times t_{m})$) (refer to Fig. 3). The fine-tuning range of the model $t_{m}$=0.25 is larger than that of the model $t_{m}$=0.10, but the performance does not exceed the latter, especially the main evaluation metrics PSNR and UIQM decreases from 24.339 and 0.943 to 23.887 and 0.894 respectively on the dataset T200. The drop in UIQM on the dataset SQUID is even more dramatic, from 0.417 to 0.291. We believe this may be due to the fact that the image-to-image diffusion model is more sensitive to high-frequency information, and that the CLIP-Classifier does not steer the generation process in the high-frequency range $t\in(0.1,0.25)$, instead interfering with the condition $y_{1}$ (refer to Fig. 3). Further comparison of the models $t_{m}$=0.50, $t_{m}$=0.75 and $t_{m}$=1.00 shows that the performance of the models is stable and the evaluation metrics are without large transition, which indicates that the CLIP-Classifier cannot distinguish the image pairs perturbed by Gaussian noise in the low frequency region $t\in(0.50,1.00)$ and has less impact on the overall generative process. This experimental result aligns with the Fig. 5 (b). Figure 12 presents the visual comparison between the proposed CLIP-UIE (the model $t_{m}$=1.00) and the other ablated versions. As shown in Fig. 12 (b), for the image-to-image special task, which is low-frequency content sharing, we are able to adapt to the downstream task by fine-tuning the model only in the range of $t\in(0,0.10)$ and the enhancement results are competitive, which is 10 times faster over the traditional strategy.
@@ V Discussion and Future works
As illustrated in Fig. 5, The score distribution of the results produced by the proposed CLIP-UIE is not clustered enough, and some enhanced images are still in the reference domain. Additionally, the score difference curve of the CLIP-Classifier is different from the normal subjective assumption—the more noise, the lower the curve—and the effective range of the classifier is discontinuous $t\in(0,0.10)\cup(0.25,0.50)$, and the image pairs perturbed by Gaussian noise cannot be distinguished when $t\in(0.10,0.25)\cup(0.50,1.00)$. This makes it difficult to precisely control the generation process of diffusion models. In future research, it would be interesting to train an effective range-stable CLIP-Classifier in order to achieve precise control of the multi-guidance diffusion model, and to make the enhancement results more natural.
As a new fine-tuning strategy, the speed is faster than the traditional fine-tuning strategy, but the overall performance is not superior to traditional methods (refer to Tab. IV). The trade-off between speed and performance is a common challenge for all deep learning-based methods. Exploring how to strike a balance speed and performance in the fine-tuning of image-to-image diffusion models is also a promising topic. Unfortunately, this fine-tuning strategy is limited to the image-to-image diffusion model, which shares low-frequency content.
@@
In this paper, we have presented a novel framework CLIP-UIE for underwater image enhancement, which combines the pre-train diffusion model and CLIP-Classifier to conduct conditional generation. In our framework, the pre-train diffusion model has learned the prior knowledge of image enhancement from the underwater degradation domain to the real in-air natural domain. The customized CLIP-Classifier inherits the prior knowledge of the in-air natural domain, which is used to counteract catastrophic forgetting and mode collapse of the diffusion model during fine-tuning, allowing the enhancement results to break through the limitations of the reference domain to a certain extent. Simultaneously, the image-to-image enhancement task requires consistent content between the input and the output images, so that the diffusion model shares low-frequency information, which means that the range of fine-tuning is concentrated in the high-frequency region. Further analysis of CLIP-Classifier shows that it also mainly acts in the high-frequency region, similar to the diffusion model. Therefore, we propose a new fine-tuning strategy that guides the pre-trained diffusion model only in the high-frequency region, e.g., the model $t_{m}$=0.10—we only fine-tune in the range $t\in(0,0.10)$—achieves competitive enhancement results, while the speed is 10 times faster than the traditional strategy. Extensive experiments show that CLIP-UIE not only generates enhanced images of superior quality beyond the reference images but also surpasses other advanced UIE methods. In the future, we aim to introduce more prior knowledge by classifiers break through the limitations of the reference domain and further enhance visual quality.

### v2

@@ IV-A Benchmarks and Evaluation Metrics
Prior Knowledge. It is well known that real reference images are difficult to obtain for the UIE task. We employ color transfer to degrade the in-air natural images of INaturalist 2021 (mini) [34] into the underwater domain to obtain synthetic paired datasets, called UIE-air dataset (see in Fig. 2 (a)). INaturalist 2021 (mini) has $10,000$ categories with $50$ images in each category, totalling $500$k images. The image-to-image diffusion model is trained from scratch on this dataset to learn the prior knowledge of mapping transitions from the real underwater degradation domain to the in-air natural domain.
Dataset. The SUIM-E [12] has only $1,635$ images, while UIEB [4] has only $890$ images, which is significantly smaller than the INaturalist dataset used to pre-train the diffusion model. Therefore, we utilize the pre-divided SUIM-E-Train set, which consists of $1,525$ paired underwater images, along with $800$ paired images randomly selected from UIEB, as the training set to fine-tune our model. The test set from SUIM-E and the remaining $90$ images from UIEB are combined to form a test set of $200$ images, called T200.
We also conduce tests on the Challenging set of UIEB, which contains $60$ tough cases for enhancement without references, called C60. Additionally, we further verify the generalization and robustness of our model on other commonly used underwater image datasets, which also lack reference images. SQUID [51] contains $57$ images taken in different locations, enabling a rigorous evaluation of restoration algorithms on underwater images. Color-Checker7 [52] covers $7$ underwater images taken with different cameras in a shallow swimming pool, which is also used to evaluate the accuracy of the proposed CLIP-UIE for color correction.
Evaluation Metrics. Since not all datasets have reference images, we divide five commonly used evaluation metrics into two groups for comprehensive quantitative evaluation. For full-reference evaluation metrics, we use Peak Signal to Noise Ratio (PSNR) and Structural Similarity (SSIM[53] to measure the consistency between the enhancement and the provided ground-truth images. For non-reference evaluation metrics, we employ three measures: Underwater Image Quality Measure (UIQM[54], which considers underwater contrast, colorfulness, and sharpness in HSV space; Underwater Color Image Quality Evaluation (UCIQE[55], which assesses chroma, contrast and saturation to quantify nonuniform color cast, blurring and low-contrast in degraded underwater images; and Cumulative Probability of Blur Detection (CPBD[56], which utilizes a probabilistic model to estimate the probability of detecting blur at each edge in the image.
@@ IV-B Implementation Details
We implement our method with the PyTorch platform. For training, we initialize all the weights of each layer with Xavier initialization [57] and employ the Adam optimizer with a learning rate of $3e$-$6$ with the scheduler of linear decay to zero. Moreover, batch size is set to $8$, and the time step of the diffusion model is set to $2000$ (namely $T=2000$). To ensure a fair comparison, images for all compared methods were resized to $256\times 256$ during both training and testing. Additionally, our data augmentation techniques include random rotation and random horizontal flips. The experiments are conducted on a consistent hardware setup comprising a single NVIDIA RTX 3090 GPU, a $2.3$GHz Intel Xeon processor, $64$GB RAM, and the Ubuntu 18.04 operating system platform. The pre-trained model we used has been trained from scratch for $300$ hours on UIE-air dataset.
@@ IV-C Compared Methods
We compare our method CLIP-UIE with seven state-of-the-art UIE approaches, including two typical traditional methods (UDCP [9] and ULAP [11]), and five deep learning-based methods (Ucolor [5], TCTL-Net [30], UIEC^2-Net [13], UDAformer [16], DM_underwater [41]). The first two UIE methods are the most representative traditional methods that have been widely adopted in recent years. The deep learning-based approaches compared here have achieved state-of-the-art (SOTA) performance over the past years. Among them, UDAformer [16] utilizes a dual attention transformer structure to establish the enhancing framework, and DM_underwater [41] represents the pioneering UIE method based on diffusion model. They demonstrate the application of the SOTA technologies in the UIE tasks. To ensure a fair and rigorous comparison, we employed the same training set from both the SUIM-E and UIEB datasets to fine-tune or further train all deep learning-based methods. Since the training phase of Ucolor  [5] requires specified image sizes, we only use $256\times 256$ image sizes for the test phase. Furthermore, other comparison methods also adopt $256\times 256$ image size in both training and testing phases. Additionally, all the comparison methods utilize the source codes released by their respective authors to generate their results and strictly adhere to identical experimental settings throughout all evaluations.
IV-D Comparison on Visual Quality of Enhancement
The quantitative comparisons on the T200, C60, SQUID and Color-Check7 datasets are reported in Table I. We mainly use PSNR and SSIM as our quantitative indices for full-reference T200 datasets, and UIQM, UCIQE and CPBD for non-reference C60, SQUID and Color-Checker7 datasets. The results in Table I show that our CLIP-UIE outperforms the other compared methods, and achieve state-of-the-art performance in terms of image quality evaluation metrics on UIE task, particularly with SSIM, UIQM and UCIQE reaching 0.936, 0.981 and 0.619 respectively. This verifies that our method can produce more natural enhanced images, which better corrects degraded colors and maps underwater images to the in-air natural domain. To better validate the superiority of our methods, Fig. 6 shows the visual results of different comparison methods on T200 dataset. The traditional methods like UDCP and ULAP usually cause color distortions and damage image detail (see the examples in Fig. 6 (b, c)), albeit with high non-reference metrics, which also occurs in other datasets. The deep learning-based methods can produce pleasing visual enhancement due to the direct use of reference images as constraints during the training phrase. However, this limits the enhanced image to a manually selected reference domain, and some methods have very similar enhancement results and do not outperform the reference (e.g. (d - h) in the second and fourth rows of Fig. 6). On the contrary, our CLIP-UIE introduces the prior knowledge of in-air natural domain through CLIP-Classifier, which can produce more natural images, breaking through the limitation imposed by reference domain to a certain extent, as shown in Fig. 6 (i).
To further verify the effectiveness and robustness of our method, we conducted experiments on C60, SQUID and Color-Checker7 datasets. (1) The visual comparisons with other methods on the C60 dataset are shown in the Fig. 7. The underwater images suffer from lower contrast and more severe color casts (see the examples in the first row of Fig. 7), which raises great challenges for the enhancement methods. By observation, the deep learning-based methods show better robustness and generalisation than the traditional methods. As a comparison, our method has better performance in color correction and detail restoration. Specially, our method achieves a 0.09 improvement over the second best method, UIEC^2-Net [13], in UIQM metrics. (2) On the SQUID dataset, our CLIP-UIE outperforms all other deep learning-based methods and achieves optimal scores in terms of underwater image quality non-reference evaluation metrics including UIQM and UCIQE, with 0.424 and 0.575, respectively. (see in Table I). As shown in Fig. 8, our method performs more natural visual enhancement for underwater images taken at different depths, which verifies the robustness of the proposed CLIP-UIE. This phenomenon can be attributed to the conditional generation is jointly controlled by the source image $y_{1}$ and in-air natural domain $y_{2}$ introduced by CLIP-Classifier (the process is shown in Fig. 3), thereby the multi-conditions guidance enhances generation quality. (3) We further conduce comparative tests on the Color-Checker7 dataset to evaluate the generalization and robustness of our model. Fig. 9 visually illustrates the difference between output images of CLIP-UIE and other methods. In the first row of Fig. 9, the difference lies in the correction of skin color, where the image generated by CLIP-UIE is closer to the real colors. In the second row of Fig. 9, the ULAP [11] produce the enhanced image with bright background lighting, but there is still a color deviation due to the introduction of excessive red components. Comparing to deep learning-based methods, the results produced by our method are brighter and more visually appealing.
@@ IV-D Comparison on Visual Quality of Enhancement
IV-E Comparison on Color Correction Performance
In order to further illustrate the performance of the proposed CLIP-UIE for color restoration of underwater images, we conduced detailed color dissimilarity comparisons on the Color-Checker7 [52] dataset, which is photographed in a swimming pool with divers holding standard color cards. Following Ancuti et al. [11], we use CIEDE2000 [52] to measure the relative differences between the corresponding color patches of the Macbeth color checker [58] and the enhancement results of these comparison state-of-the-art methods. For better illustration, a standard color card is appended to the upper left corner of the first column of Fig. 9. The results, as presented in Table II, indicate that our CLIP-UIE can improve the performance in recovering ground-truth colors, but not very noticeably. However, the color performance of the image T8000 is even worse than the input. We believe that this is due to the image encoder $\Phi(\cdot)$ compresses images into high-dimensional features, and CLIP-Classifier pays more attention to global features and ignores the color restoration of local color blocks during generation process. As shown in the first row of Fig. 9, the result of our method presents better brightness and visual perception, but does not achieve a significant improvement in color correction.
@@ IV-F Ablation Study
Effectiveness of CLIP-Classifier. To Investigate the impact of the CLIP-Classifier on our method, we fine-tune our conditional diffusion model on the same training data without CLIP-Classifier (w/o-CLIP). The scores of the main evaluation metrics on T200 datasets are given in Table III. The results indicate that CLIP-Classifier guides the diffusion model toward the in-air natural domain, leading to a decrease in the full-reference metrics and an increase in the non-reference metrics. Besides, we use CLIP-Classifier to evaluate the similarity (CLIP score) between the in-air natural domain and enhanced results, and the corresponding point-box plot and kernel density estimation (KDE) curve are shown in Fig. 10. The distributions of the w/o-CLIP and the Reference are very similar (see KDE in Fig. 10), which indicates that its enhancement performance is constrained by the reference domain. On the contrary, due to the introduction of the prior knowledge of in-air natural domain through CLIP-Classifier, our CLIP-UIE breaks through this limitation imposed by reference domain to a certain extent, keeping the distribution of the CLIP-UIE far away from the reference domain and close to the in-air natural domain, reducing subjective preferences.
Visual comparison illustrated in Fig. 11 can also significantly support the above points. Compared with its ablated versions, the proposed CLIP-UIE achieves more natural enhancement results with the guidance of CLIP-Classifier that inherits the prior knowledge of the in-air natural domain. The combination of the reference domain and the in-air natural domain drives the enhancement results closer to real-world natural images (see Fig. 3 and Fig. 11 (c, d)).
@@
Fine-Tuning Strategy. In addition to the observation provided in Section III-C, to further validate the new fine-tuning strategy, we provide the quantitative comparison in Table IV. The model $t_{m}$ represents the fact that we only fine-tune the pretrained diffusion model in the range $t\in(0,t_{m})$ (the training time step $n\in(0,2000\times t_{m})$) (refer to Fig. 3). As showm in Table IV, the fine-tuning range of the model $t_{m}$=0.25 is larger than that of the model $t_{m}$=0.10, but the performance does not exceed the latter, especially the main evaluation metrics PSNR and UIQM decreases from 24.339 and 0.943 to 23.887 and 0.894 respectively on the dataset T200. The drop in UIQM on the dataset SQUID is even more dramatic, from 0.417 to 0.291. We believe this may be due to the fact that the image-to-image diffusion model is more sensitive to high-frequency information. The CLIP-Classifier not only does not steer the generation process in the high-frequency range $t\in(0.1,0.25)$, but also interferes with the condition $y_{1}$ (refer to Fig. 3). Further comparison of the models $t_{m}$=0.50, $t_{m}$=0.75 and $t_{m}$=1.00 shows that the performance of the models is stable and there is no drastic change in the evaluation metrics. This indicates that the CLIP-Classifier cannot distinguish the image pairs perturbed by Gaussian noise in the low-frequency region $t\in(0.50,1.00)$ and has less impact on the overall generative process. This experimental result aligns with the Fig. 5 (b). Fig. 12 presents the visual comparison between the proposed CLIP-UIE (the model $t_{m}$=1.00) and the other ablated versions. As shown in Fig. 12 (b), for the image-to-image special task, which shares low-frequency semantic information, we are able to adapt to the downstream task by fine-tuning the model only in the range of $t\in(0,0.10)$ and the enhancement results are competitive, which is 10 times faster over the traditional strategy.
@@ V Discussion and Future works
As illustrated in Fig. 10, the score distribution of the results produced by the proposed CLIP-UIE is not clustered enough, and some enhanced images are still in the reference domain. Additionally, the score difference curve of the CLIP-Classifier is different from the normal subjective assumption—the more noise, the lower the curve—and the effective range of the classifier is discontinuous $t\in(0,0.10)\cup(0.25,0.50)$, and the image pairs perturbed by Gaussian noise cannot be distinguished when $t\in(0.10,0.25)\cup(0.50,1.00)$. This makes it difficult to precisely control the generation process of diffusion models. In future research, it would be interesting to train an effective range-stable CLIP-Classifier to achieve precise control of the multi-guidance diffusion model, and to make the enhancement results more natural.
As a new fine-tuning strategy, the speed is faster than the traditional fine-tuning strategy, but the overall performance is not superior to traditional methods (refer to Table IV). The trade-off between speed and performance is a common challenge for all deep learning-based methods. Exploring how to strike a balance between speed and performance in the fine-tuning of image-to-image diffusion models is also a promising topic. Unfortunately, this fine-tuning strategy is limited to the image-to-image diffusion model, which shares low-frequency semantic information.
@@
In this paper, we have presented a novel framework called CLIP-UIE for underwater image enhancement, which combines the pre-trained diffusion model and CLIP-Classifier to conduct conditional generation. In our framework, the pre-trained diffusion model has learned the prior knowledge of mapping transitions from the underwater degradation domain to the real in-air natural domain. The customized CLIP-Classifier inherits the prior knowledge of the in-air natural domain, which is used to counteract catastrophic forgetting and mode collapse of the diffusion model during fine-tuning, allowing the enhancement results to break through the limitations of the reference domain to a certain extent. Simultaneously, the image-to-image enhancement task requires consistent content between the input and the output images, so that the diffusion model shares low-frequency semantic information, which means that the range of fine-tuning is concentrated in the high-frequency region. Further analysis of CLIP-Classifier shows that it also mainly acts in the high-frequency region, similar to the image-to-image diffusion model. Therefore, we propose a new fine-tuning strategy that guides the pre-trained diffusion model only in the high-frequency region, e.g., the model $t_{m}$=0.10—we only fine-tune in the range $t\in(0,0.10)$—achieves competitive enhancement results, while the speed is 10 times faster than the traditional strategy. Extensive experiments show that CLIP-UIE not only generates enhanced images of superior quality beyond the reference images, but also surpasses other SOTA UIE methods. In the future, we aim to introduce more prior knowledge by classifiers, and further enhance visual quality.